# Newton’s method for nonlinear mappings into vector bundles

### v1

@@ Abstract
We consider Newton’s method for finding zeros of mappings from a manifold $\mathcal{X}$ into a vector bundle $\mathcal{E}$. In this setting a connection on $\mathcal{E}$ is required to render the Newton equation well defined, and a retraction on $\mathcal{X}$ is needed to compute a Newton update. We discuss local convergence in terms of suitable differentiability concepts, using a Banach space variant of a Riemannian distance. We also carry over an affine covariant damping strategy to our setting. Finally, we discuss two simple applications of our approach, namely, finding fixed points of vector fields and stationary points of functionals.
@@ 1. Introduction
Newton’s method is a central algorithm for the solution of nonlinear problems, but also, in its variants, a theoretical tool, e.g., in the proof of the implicit function theorem [21]. Most of the literature consider Newton’s method for finding zeros $F(x)=0$ for a differentiable mapping $F:X\to Y$, where $X$ and $Y$ are linear, normed, not necessarily finite dimensional spaces, in most cases Banach spaces. In this setting, a Newton step is defined straightforwardly as the solution $\delta x\in X$ of the linear operator equation and a simple additive update:
@@
Extensions of Newton’s method to problems $F:\mathcal{X}\to Y$, where $\mathcal{X}$ is a Riemannian manifold and $Y$ still a linear space with the help of so called retractions $R_{x}:T_{x}\mathcal{X}\to\mathcal{X}$, are relatively straightforward, replacing the simple update by $x_{+}=R_{x}(\delta x)$. But also, Newton’s method has been used to find stationary points of twice differentiable functions $f:\mathcal{X}\to\mathbb{R}$ [1], and finally to find zeros of vector fields $v\in\Gamma(\mathcal{X})$ on $\mathcal{X}$ [2, 6]. However, much richer classes of problems are conceivable in a manifold setting, such as constrained optimization [3, 15, 22, 23], non-symmetric variational problems, problems of stationary action [16], or other saddle point problems, to name just a few.
@@
Specifically, we want to establish Newton’s method for a mapping $F:\mathcal{X}\to\mathcal{E}$, where $\mathcal{X}$ is a differentiable manifold and $p:\mathcal{E}\to\mathcal{M}$ is a vector bundle with fibres $E_{y}$, $y\in\mathcal{M}$. This formulation still allows the search for zeros, i.e. $F(x)=0_{p(F(x))}\in E_{p(F(x))}$, and it covers many of the above mentioned settings. The derivative of $F$ is a mapping $F^{\prime}(x):T_{x}\mathcal{X}\to T_{F(x)}\mathcal{E}$, which means the Newton equation cannot be formulated in the usual way, since the codomains of $F$ and $F^{\prime}(x)$, i.e., $\mathcal{E}$ and $T_{F(x)}\mathcal{E}$, do not coincide. It is thus necessary to introduce some additional geometric structure. It turns out that we need a connection on $\mathcal{E}$, which induces for every $e\in\mathcal{E}$ a linear projection $Q_{e}:T_{e}\mathcal{E}\to E_{p(e)}$ onto the corresponding fibre. This allows us to formulate the Newton equation in a well defined way:
@@
If $Q_{F(x)}\circ F^{\prime}(x)$ is invertible, the Newton direction $\delta x\in T_{x}\mathcal{X}$ can be computed, and thus the Newton step via a retraction $x_{+}=R_{x}(\delta x)$.
To observe convergence of Newton’s method we need a metric $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ on $\mathcal{X}$. Since Newton’s method in the linear setting is not restricted to Hilbert spaces and has important applications in Banach spaces (cf. e.g. [25] or [7]), we will introduce metrics similar to, but more general than Riemannian distances. Finally, to compare images of successive iterates on different fibres, e.g., $F(x_{+})$ and $F(x)$, vector back-transports of the form $V^{-1}_{y}:\mathcal{E}\to E_{y}$ are needed on $\mathcal{E}$ (we call them back-transports, because they work in the opposite direction than classical vector transports).
After setting up this framework, we will establish basic local convergence results, based on minimal assumptions on the smoothness of $F$, using a geometric version of Newton differentiability. Besides giving a classical a-priori result, we also establish a theorem, which relates local convergence to a quantity that is purely defined in terms of the space of iterates $\mathcal{X}$, and which can be estimated a-posteriori. This opens the door to define an affine covariant damping strategy in the spirit of Deuflhard [9], which is based on following the so called Newton path. While carrying over this strategy from linear spaces to our nonlinear setting, we will observe that strict differentiability of $F$ is required to make this strategy viable. In addition, it will turn out that the employed connection and the vector back-transport have to satisfy a consistency condition to ensure that useful steps are taken. Finally, we will illustrate our results for the two simplest applications, namely the search for fixed points of vector fields and stationary points of functions on $\mathcal{X}$.
@@ 2. Preliminaries
For a manifold $\mathcal{X}$ we define tangent spaces by using charts, see, e.g., [13, II §2]. Let $x\in\mathcal{X}$ and consider the set
@@
The relation $(\phi_{i},v_{i})\sim(\phi_{j},v_{j})\ :\Leftrightarrow\ t_{ji}^{\prime}(\phi_{i}(x))v_{i}=v_{j}$ defines an equivalence relation. The corresponding equivalence class is called a tangent vector $v$ of $\mathcal{X}$ at $x$. The set of these equivalence classes at a point $x\in\mathcal{X}$ is a vector space, called the tangent space $T_{x}\mathcal{X}$ of $\mathcal{X}$ at $x$. Every differentiable map $F:\mathcal{X}\rightarrow\mathcal{M}$ between two $C^{1}$-manifolds $\mathcal{X}$ and $\mathcal{M}$ induces a linear map $F^{\prime}(x_{0}):T_{x_{0}}\mathcal{X}\rightarrow T_{F(x_{0})}\mathcal{M}$, termed tangent map or derivative of $F$ at $x_{0}\in\mathcal{X}$ [13].
@@
Consider a vector bundle $p:\mathcal{E}\rightarrow\mathcal{M}$. Its projection $p$ is a smooth surjective map, which assigns to $e\in\mathcal{E}$ its base point $y=p(e)$ in the base manifold $\mathcal{M}$ [13, III §1]. The total space $\mathcal{E}$ is a manifold with special structure: for each $y\in\mathcal{M}$ the preimage $E_{y}:=p^{-1}(y)$ is a Banachable space (i.e. a complete topological vector space whose topology is induced by some norm, rendering $E_{y}$ complete), called fibre over $y$. The zero element in $E_{y}$ is denoted by $0_{y}$. A mapping $v:\mathcal{M}\to\mathcal{E}$, such that $p(v(y))=y$ for all $y\in\mathcal{M}$ is called section of $\mathcal{E}$. The set of all sections is denoted by $\Gamma(\mathcal{E})$. The most prominent example for a vector bundle is the tangent bundle $\pi:T\mathcal{X}\to\mathcal{X}$ of a manifold $\mathcal{X}$ with fibres $T_{x}\mathcal{X}$. A section of $T\mathcal{X}$ is called a vector field.
Vector bundles can be described via general local charts, but usually, local trivializations are used instead, which reflect their structure better. For a Banach space $(\mathbb{E},\|\cdot\|_{\mathbb{E}})$ and an open set $U\subset\mathcal{M}$, a local trivialization is a diffeomorphism $\tau:p^{-1}(U)\to U\times\mathbb{E}$, such that $\tau_{y}:=\tau|_{E_{y}}\in L(E_{y},\mathbb{E})$ is a linear isomorphism. Thus, any element of $e\in\mathcal{E}$ can be represented as a pair $(y,e_{\tau})\in\mathcal{M}\times\mathbb{E}$ with $y=p(e)$ and $e_{\tau}$ depending on the trivialization. For two trivializations $\tau$ and $\tilde{\tau}$ around $y\in\mathcal{M}$ we can define smooth transition mappings $A:y\mapsto\tilde{\tau}_{y}\tau_{y}^{-1}$ with $A(y)\in L(\mathbb{E},\mathbb{E})$, and obtain $e_{\tilde{\tau}}=A(y)e_{\tau}$. If $\phi:U\to\mathbb{M}$ is a chart of $\mathcal{M}$, then we can construct a vector bundle chart:
$\Phi_{\mathcal{E}}:p^{-1}(U)$ $\to\mathbb{M}\times\mathbb{E},\quad e\mapsto(\phi(p(e)),\tau_{p(e)}e),$
@@
Let $(U,\phi_{i})$ be a chart of $\mathcal{X}$ at $x$. Then every tangent vector $v\in T_{x}\mathcal{X}$ has a chart representation $v_{i}\in\mathbb{X}$, so $\tau_{i}(v):=v_{i}$ is a trivialization. Since tangent vectors transform by $t_{ji}^{\prime}(\phi_{i}(x))v_{i}=v_{j}$ under change of charts, the transition mapping is given by $A(x)=t_{ji}^{\prime}(\phi_{i}(x))$. A vector bundle chart for $T\mathcal{X}$ is then given by:
$\Phi_{T\mathcal{X}}:\pi^{-1}(U)$ $\to\mathbb{X}\times\mathbb{X},\quad(x,v)\mapsto(\phi_{i}(x),v_{i}).$
Similarly, each element $\delta e\in T_{e}\mathcal{E}$ can be represented in local trivializations as a pair of elements $(\delta y,\delta e_{\tau})\in T_{y}\mathcal{M}\times\mathbb{E}$, where the tangential part $\delta y=p^{\prime}(e)\delta e$ is given canonically. The representation $\delta e_{\tau}$ of the fibre part, however, depends crucially on the chosen trivialization, and so there is no natural splitting of $\delta e$ into a tangential part and a fibre part. In fact, if $e_{\tilde{\tau}}=A(y)e_{\tau}$, then $\delta e_{\tilde{\tau}}=A(y)\delta e_{\tau}+(A^{\prime}(y)\delta y)e_{\tau}$ by the product rule.
@@
Consider two vector bundles $p_{1}:\mathcal{E}_{1}\to\mathcal{M}_{1},p_{2}:\mathcal{E}_{2}\to\mathcal{M}_{2}$, and denote by $L(E_{1,y},E_{2,z})$ the set of continuous linear mappings $E_{1,y}\to E_{2,z}$ for $y\in\mathcal{M}_{1}$ and $z\in\mathcal{M}_{2}$. Given a mapping $f:\mathcal{M}_{1}\to\mathcal{M}_{2}$ we denote by $f^{*}\mathcal{E}_{2}$ the pullback bundle of $\mathcal{E}_{2}$ via $f$, i.e. $(f^{*}E_{2})_{y}:=E_{2,f(y)}$. Collecting all mappings in $L(E_{1,y},E_{2,f(y)})$ for all $y\in\mathcal{M}_{1}$ in a set $\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2})$, we observe that $p_{\mathcal{L}}:\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2})\to\mathcal{M}_{1}$ is a vector bundle itself with fibres $L(E_{1,y},E_{2,f(y)})$. In trivializations an element $H$ of $\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2})$ is represented by a pair $(y,H_{\tau})$ where $y\in\mathcal{M}_{1}$ and $H_{\tau}\in L(\mathbb{E}_{1},\mathbb{E}_{2})$.
A fibrewise linear mapping, also called vector bundle morphism, see e.g., [13, III §1], is a section $S\in\Gamma(\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2}))$, i.e, a mapping
$S:\mathcal{M}_{1}\to\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2})\text{ with }p_{\mathcal{L}}\circ S=Id_{\mathcal{M}_{1}}.$
For example, if $F:\mathcal{X}\to\mathcal{M}$ is differentiable, then the derivative $S:=F^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{X},F^{*}T\mathcal{M}))$ is a fibrewise linear mapping, but also vector transports, defined below, are fibrewise linear mappings. $S$ is called locally bounded, respectively differentiable, if its representation in local trivializations, i.e.
$S_{\tau}:=\tau_{\mathcal{E}_{2}}\circ S\circ\tau_{\mathcal{E}_{1}}^{-1}$ (1)
has the same property. If $S$ is differentiable, then its derivative is a section $S^{\prime}\in\Gamma(T\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2}))$ of the tangent bundle $p_{\mathcal{L}}^{\prime}:T\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2})\to T\mathcal{M}_{1}$, i.e., a mapping
$S^{\prime}:T\mathcal{M}_{1}\to T\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2}),\text{ such that }p_{\mathcal{L}}^{\prime}\circ S^{\prime}=Id_{T\mathcal{M}_{1}}.$
From $S\in\Gamma(\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2}))$ we can construct a mapping ${\langle{S}\rangle}\in C^{1}(\mathcal{E}_{1},\mathcal{E}_{2})$ as follows:
@@
$\varphi:\Gamma(\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2}))\to C^{1}(\mathcal{E}_{1},\mathcal{E}_{2}),\quad S\mapsto{\langle{S}\rangle}.$
Comparison of $S^{\prime}\in\Gamma(T\mathcal{L}(\mathcal{E}_{1},f^{*}\mathcal{E}_{2}))$ with ${\langle{S}\rangle}^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{E}_{1},S^{*}T\mathcal{E}_{2}))$ in trivializations yields by the product rule
${\langle{S}\rangle}_{\tau}^{\prime}(e)(\delta y,\delta e_{\tau})=S_{\tau}(p_{1}(e))\delta e_{\tau}+(S_{\tau}^{\prime}(p_{1}(e))\delta y)e_{\tau}\quad\forall(\delta y,\delta e_{\tau})\in T_{p_{1}(e)}\mathcal{M}_{1}\times\mathbb{E}_{1}.$ (2)
@@
For a vector bundle $p:\mathcal{E}\rightarrow\mathcal{M}$ and $y\in\mathcal{M}$ we define a vector transport as a section $V_{y}\in\Gamma(\mathcal{L}(\mathcal{M}\times E_{y},\mathcal{E}))$, i.e.
$V_{y}:\mathcal{M}\to\mathcal{L}(\mathcal{M}\times E_{y},\mathcal{E}),\ \text{with }p_{\mathcal{L}}(V_{y}(\hat{y}))=\hat{y}\; \forall\hat{y}\in\mathcal{M},$
with the properties that $V_{y}(y)=Id_{E_{y}}$ and $V_{y}(\hat{y})$ is invertible for all $\hat{y}\in\mathcal{M}$.
@@
$V_{y}^{-1}:\mathcal{M}\to\mathcal{L}(\mathcal{E},E_{y}),\ \text{with }p_{\mathcal{L}}(V^{-1}_{y}(\hat{y}))=\hat{y}\; \forall\hat{y}\in\mathcal{M}$
with the property that $V^{-1}_{y}(y)=Id_{E_{y}}$ and $V^{-1}_{y}(\hat{y})$ is invertible for all $\hat{y}\in\mathcal{M}$.
By restricting $\mathcal{M}$ appropriately, vector transports can also be defined locally. Given a vector transport $V_{y}\in\Gamma(\mathcal{L}(\mathcal{M}\times E_{y},\mathcal{E})$, we can derive a vector back-transport $V_{y}^{-1}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ by taking pointwise inverses $V_{y}^{-1}(\hat{y}):=(V_{y}(\hat{y}))^{-1}$, $\hat{y}\in\mathcal{M}$, and vice versa.

### v2

@@ Abstract
We consider Newton’s method for finding zeros of mappings from a manifold $\mathcal{X}$ into a vector bundle $\mathcal{E}$. In this setting a connection on $\mathcal{E}$ is required to render the Newton equation well defined, and a retraction on $\mathcal{X}$ is needed to compute a Newton update. We discuss local convergence in terms of suitable differentiability concepts, using a Banach space variant of a Riemannian distance. We also carry over an affine covariant damping strategy to our setting. Finally, we will discuss some applications of our approach, namely, finding fixed points of vector fields, variational problems on manifolds and finding critical points of functionals.
@@ 1. Introduction
Newton’s method is a central algorithm for the solution of nonlinear problems, but also, in its variants, a theoretical tool, e.g., in the proof of the implicit function theorem. Most of the literature consider Newton’s method for finding zeros $F(x)=0$ for a differentiable mapping $F:X\to Y$, where $X$ and $Y$ are linear, normed, not necessarily finite dimensional spaces, in most cases Banach spaces. In this setting, a Newton step is defined as the solution $\delta x\in X$ of a linear operator equation together with an additive update:
@@
Extensions of Newton’s method to problems $F:\mathcal{X}\to Y$, where $\mathcal{X}$ is a Riemannian manifold and $Y$ still a linear space with the help of so called retractions $R_{x}:T_{x}\mathcal{X}\to\mathcal{X}$, are relatively straightforward, replacing the additive update by $x_{+}=R_{x}(\delta x)$. But also, Newton’s method has been used to find stationary points of twice differentiable functions $f:\mathcal{X}\to\mathbb{R}$ [1], and finally to find zeros of vector fields $v\in\Gamma(\mathcal{X})$ on $\mathcal{X}$ [2, 6]. However, much richer classes of problems are conceivable in a manifold setting, such as constrained optimization [3, 16, 24, 25], non-symmetric variational problems, problems of stationary action [17], or other saddle point problems, to name just a few.
@@
Specifically, we want to establish Newton’s method for a mapping $F:\mathcal{X}\to\mathcal{E}$, where $\mathcal{X}$ is a differentiable manifold and $p:\mathcal{E}\to\mathcal{Y}$ is a vector bundle with fibres $E_{y}$, $y\in\mathcal{Y}$. This setting still allows to formulate root-finding problems, which now read $F(x)=0_{p(F(x))}\in E_{p(F(x))}$, but it is general enough to cover the above mentioned problem classes.
In contrast to the classical case the linear space $E_{p(F(x))}$ in which the residuals are evaluated is not constant, but depends on $x$ via $p\circ F$. Moreover, the derivative of $F:\mathcal{X}\to\mathcal{E}$ is now a mapping $F^{\prime}(x):T_{x}\mathcal{X}\to T_{F(x)}\mathcal{E}$, which means that the Newton equation cannot be formulated in the usual way, since the codomains of $F$ and $F^{\prime}(x)$, i.e., $\mathcal{E}$ and $T_{F(x)}\mathcal{E}$, do not coincide. It is thus necessary to introduce some additional geometric structure, namely a connection on $\mathcal{E}$. A conncetion induces for every $e\in\mathcal{E}$ a linear mapping $Q_{e}:T_{e}\mathcal{E}\to E_{p(e)}$, which plays the role of a projection onto the corresponding fibre. This allows us to formulate the Newton equation in a well defined way:
@@
If $Q_{F(x)}\circ F^{\prime}(x)$ is invertible, the Newton direction $\delta x\in T_{x}\mathcal{X}$ can be computed, and then the Newton step can be defined via a retraction $x_{+}=R_{x}(\delta x)$.
To observe convergence of Newton’s method we need a metric $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ on $\mathcal{X}$. Since Newton’s method in the linear setting is not restricted to Hilbert spaces and has important applications in Banach spaces (cf. e.g. [27] or [7]), we will introduce metrics similar to, but more general than Riemannian distances. Finally, to compare images of successive iterates on different fibres, e.g., $F(x_{+})\in E_{p(F(x_{+}))}$ and $F(x)\in E_{p(F(x))}$, vector back-transports of the form $V^{-1}_{y}:\mathcal{E}\to E_{y}$ are needed on $\mathcal{E}$ (we call them back-transports, because they work in the opposite direction than classical vector transports).
After setting up this framework, we will establish basic local convergence results, based on minimal assumptions on the smoothness of $F$, using a geometric version of Newton differentiability. Besides giving a classical a-priori result, we also establish a theorem, which relates local convergence to a quantity that is purely defined in terms of the space of iterates $\mathcal{X}$, and which can be estimated a-posteriori. This opens the door to define an affine covariant damping strategy in the spirit of Deuflhard [9], which is based on following the so called Newton path. While carrying over this algorithm from linear spaces to our nonlinear setting, we will observe that strict differentiability of $F$ is required together with a consistency condition of the employed connection and the vector back-transport to make this strategy viable. Finally, we will indicate, how our results can be applied to the search for fixed points of vector fields, to variational problems on manifolds and to finding critical points of functionals. We will also discuss these applications in the important special case that $\mathcal{E}$ is an embedded vector bundle.
@@ 2. Preliminaries
For a manifold $\mathcal{X}$ we define tangent spaces by using charts, see, e.g., [14, II §2]. Let $x\in\mathcal{X}$ and consider the set
@@
The relation $(\phi_{i},v_{i})\sim_{x}(\phi_{j},v_{j})\ :\Leftrightarrow\ t_{ji}^{\prime}(\phi_{i}(x))v_{i}=v_{j}$ defines an equivalence relation. The corresponding equivalence class is called a tangent vector $v$ of $\mathcal{X}$ at $x$. The set of these equivalence classes at a point $x\in\mathcal{X}$ is a vector space, called the tangent space $T_{x}\mathcal{X}$ of $\mathcal{X}$ at $x$. Every differentiable map $F:\mathcal{X}\rightarrow\mathcal{Y}$ between two $C^{1}$-manifolds $\mathcal{X}$ and $\mathcal{Y}$ induces a linear map $F^{\prime}(x):T_{x}\mathcal{X}\rightarrow T_{F(x)}\mathcal{Y}$, termed tangent map or derivative of $F$ at any $x\in\mathcal{X}$ [14].
@@
Consider a vector bundle $p:\mathcal{E}\rightarrow\mathcal{Y}$, which will be assumed to be of class $C^{1}$, unless otherwise noted. Its projection $p$ is a surjective $C^{1}$-map, which assigns to $e\in\mathcal{E}$ its base point $y=p(e)$ in the base manifold $\mathcal{Y}$ [14, III §1]. The total space $\mathcal{E}$ is a manifold with special structure: for each $y\in\mathcal{Y}$ the preimage $E_{y}:=p^{-1}(y)$ is a Banachable space (i.e. a complete topological vector space whose topology is induced by some norm, rendering $E_{y}$ complete), called fibre over $y$. The zero element in $E_{y}$ is denoted by $0_{y}$. A mapping $v:\mathcal{Y}\to\mathcal{E}$, such that $p(v(y))=y$ for all $y\in\mathcal{Y}$ is called section of $\mathcal{E}$. The set of all sections is denoted by $\Gamma(\mathcal{E})$. The most prominent example for a vector bundle is the tangent bundle $\pi:T\mathcal{X}\to\mathcal{X}$ of a manifold $\mathcal{X}$ with fibres $T_{x}\mathcal{X}$. A section of $T\mathcal{X}$ is called a vector field.
Vector bundles can be described via general local charts, but usually, local trivializations are used instead, which reflect their structure better. For a Banach space $(\mathbb{E},\|\cdot\|_{\mathbb{E}})$ and an open set $U\subset\mathcal{Y}$, a local trivialization is a diffeomorphism $\tau:p^{-1}(U)\to U\times\mathbb{E}$, such that $\tau_{y}:=\tau|_{E_{y}}\in L(E_{y},\mathbb{E})$ is a linear isomorphism. Thus, any element of $e\in\mathcal{E}$ can be represented as a pair $(y,e_{\tau})\in\mathcal{Y}\times\mathbb{E}$ with $y=p(e)$ and $e_{\tau}$ depending on the trivialization. For two trivializations $\tau$ and $\tilde{\tau}$ around $y\in\mathcal{Y}$ we can define smooth transition mappings $A:y\mapsto\tilde{\tau}_{y}\tau_{y}^{-1}$ with $A(y)\in L(\mathbb{E},\mathbb{E})$, and obtain $e_{\tilde{\tau}}=A(y)e_{\tau}$. If $\phi:U\to\mathbb{Y}$ is a chart of $\mathcal{Y}$, then we can construct a vector bundle chart:
$\Phi_{\mathcal{E}}:p^{-1}(U)$ $\to\mathbb{Y}\times\mathbb{E},\quad e\mapsto(\phi(p(e)),\tau_{p(e)}e),$
@@
Let $(U_{i},\phi_{i})$ be a chart of $\mathcal{X}$ at $x$. Then every tangent vector $v\in T_{x}\mathcal{X}$ has a chart representation $v_{i}\in\mathbb{X}$, so $\tau_{i}(v):=v_{i}$ is a trivialization. Since tangent vectors transform by $t_{ji}^{\prime}(\phi_{i}(x))v_{i}=v_{j}$ under change of charts, the transition mapping is given by $A(x)=t_{ji}^{\prime}(\phi_{i}(x))$. A vector bundle chart for $T\mathcal{X}$ is then given by:
$\Phi_{T\mathcal{X}}:\pi^{-1}(U_{i})$ $\to\mathbb{X}\times\mathbb{X},\quad(x,v)\mapsto(\phi_{i}(x),v_{i}).$
Similarly, each element $\delta e\in T_{e}\mathcal{E}$ can be represented in local trivializations as a pair of elements $(\delta y,\delta e_{\tau})\in T_{y}\mathcal{Y}\times\mathbb{E}$, where the tangential part $\delta y=p^{\prime}(e)\delta e$ is given canonically. The representation $\delta e_{\tau}$ of the fibre part, however, depends crucially on the chosen trivialization, and so there is no natural splitting of $\delta e$ into a tangential part and a fibre part. In fact, if $e_{\tilde{\tau}}=A(y)e_{\tau}$, then $\delta e_{\tilde{\tau}}=A(y)\delta e_{\tau}+(A^{\prime}(y)\delta y)e_{\tau}$ by the product rule.
Given a mapping $f:\mathcal{X}\to\mathcal{Y}$ and a vector bundle $p:\mathcal{E}\to\mathcal{Y}$ we denote by $f^{*}p:f^{*}\mathcal{E}\to\mathcal{X}$ the pullback bundle of $\mathcal{E}$ via $f$, i.e. $(f^{*}E)_{x}:=E_{f(x)}$, or $f\circ(f^{*}p)=p$.
@@
Consider two vector bundles $p_{1}:\mathcal{E}_{1}\to\mathcal{Y}_{1},\;p_{2}:\mathcal{E}_{2}\to\mathcal{Y}_{2}$, and denote by $L(E_{1,y},E_{2,z})$ the set of continuous linear mappings $E_{1,y}\to E_{2,z}$ for $y\in\mathcal{Y}_{1}$ and $z\in\mathcal{Y}_{2}$. Given a mapping $s:\mathcal{Y}_{1}\to\mathcal{Y}_{2}$
we can collect all linear mappings in $L(E_{1,y},E_{2,s(y)})$ for all $y\in\mathcal{Y}_{1}$ in a set $\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2})$. We observe that $p_{\mathcal{L}}:\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2})\to\mathcal{Y}_{1}$ is a vector bundle itself with fibres $L(E_{1,y},E_{2,s(y)})$. In trivializations an element $H$ of $\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2})$ is represented by a pair $(y,H_{\tau})$ where $y\in\mathcal{Y}_{1}$ and $H_{\tau}\in L(\mathbb{E}_{1},\mathbb{E}_{2})$. If the co-domain is a fixed fibre, i.e, $E_{z}\cong\{z\}\times E_{z}$, we use the simplified notation $\mathcal{L}(\mathcal{E}_{1},E_{z})$ instead of the cumbersome $\mathcal{L}(\mathcal{E}_{1},s_{z}^{*}(\{z\}\times E_{z}))$, where $s_{z}:\mathcal{Y}_{1}\to\{z\}$ is the constant map.
We define a fibrewise linear mapping, also known as vector bundle morphism, see, e.g., [14, III §1], as a section $S\in\Gamma(\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2}))$, i.e, a mapping
$S:\mathcal{Y}_{1}\to\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2})\text{ with }p_{\mathcal{L}}\circ S=Id_{\mathcal{Y}_{1}}.$
$S$ is called locally bounded, respectively differentiable, if its representation in local trivializations
$S_{\tau}:U\to L(\mathbb{E}_{1},\mathbb{E}_{2}),\quad S_{\tau}(y)=\tau_{\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2})}(S):=\tau_{\mathcal{E}_{2},s(y)}\circ S(y)\circ\tau_{\mathcal{E}_{1},y}^{-1}\in L(\mathbb{E}_{1},\mathbb{E}_{2})$ (1)
has the respective property.
If $S$ is differentiable, then its derivative is a section $S^{\prime}\in\Gamma(T\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2}))$ of the tangent bundle $p_{\mathcal{L}}^{\prime}:T\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2})\to T\mathcal{Y}_{1}$, i.e., a mapping
$S^{\prime}:T\mathcal{Y}_{1}\to T\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2}),\text{ such that }p_{\mathcal{L}}^{\prime}\circ S^{\prime}=Id_{T\mathcal{Y}_{1}}.$
From $S\in\Gamma(\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2}))$ we can construct a mapping ${\langle{S}\rangle}\in C^{1}(\mathcal{E}_{1},\mathcal{E}_{2})$ as follows:
@@
$\varphi:\Gamma(\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2}))\to C^{1}(\mathcal{E}_{1},\mathcal{E}_{2}),\quad S\mapsto{\langle{S}\rangle}.$
Comparison of the derivatives $S^{\prime}\in\Gamma(T\mathcal{L}(\mathcal{E}_{1},s^{*}\mathcal{E}_{2}))$ and ${\langle{S}\rangle}^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{E}_{1},S^{*}T\mathcal{E}_{2}))$ in trivializations yields by the product rule
${\langle{S}\rangle}_{\tau}^{\prime}(y,e_{\tau})(\delta y,\delta e_{\tau})=S_{\tau}(y)\delta e_{\tau}+(S_{\tau}^{\prime}(y)\delta y)e_{\tau}\quad\forall(\delta y,\delta e_{\tau})\in T_{y}\mathcal{Y}_{1}\times\mathbb{E}_{1}.$ (2)
For example, if $F:\mathcal{X}\to\mathcal{Y}$ is twice differentiable, then the derivative $S:=F^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{X},F^{*}T\mathcal{Y}))$ is a differentiable fibrewise linear mapping with $\langle F^{\prime}\rangle\in C^{1}(T\mathcal{X},T\mathcal{Y})$.
@@
For a vector bundle $p:\mathcal{E}\rightarrow\mathcal{Y}$ and $y\in\mathcal{Y}$ we define a vector transport as a section $V_{y}\in\Gamma(\mathcal{L}(\mathcal{Y}\times E_{y},\mathcal{E}))$, i.e.
$V_{y}:\mathcal{Y}\to\mathcal{L}(\mathcal{Y}\times E_{y},\mathcal{E}),\ \text{with }p_{\mathcal{L}}(V_{y}(\hat{y}))=\hat{y}\; \forall\hat{y}\in\mathcal{Y},$
with the properties that $V_{y}(y)=Id_{E_{y}}$ and $V_{y}(\hat{y})$ is invertible for all $\hat{y}\in\mathcal{Y}$.
@@
$V_{y}^{-1}:\mathcal{Y}\to\mathcal{L}(\mathcal{E},E_{y}),\ \text{with }p_{\mathcal{L}}(V^{-1}_{y}(\hat{y}))=\hat{y}\; \forall\hat{y}\in\mathcal{Y}$
with the property that $V^{-1}_{y}(y)=Id_{E_{y}}$ and $V^{-1}_{y}(\hat{y})$ is invertible for all $\hat{y}\in\mathcal{Y}$.
By restricting $\mathcal{Y}$ appropriately, vector transports can also be defined locally. Given a vector transport $V_{y}\in\Gamma(\mathcal{L}(\mathcal{Y}\times E_{y},\mathcal{E}))$, we can derive a vector back-transport $V_{y}^{-1}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ by taking fibrewise inverses $V_{y}^{-1}(\hat{y}):=(V_{y}(\hat{y}))^{-1}$, $\hat{y}\in\mathcal{Y}$, and vice versa. Often, vector transports are given as mappings $V(\cdot,\cdot):\mathcal{Y}\times\mathcal{Y}\to\mathcal{L}(\mathcal{E},\mathcal{E})$ with $V(y,\hat{y})\in L(E_{y},E_{\hat{y}})$. These admit interpretations as vector transports or as vector back-transports as follows: $V_{y}(\hat{y})=V(y,\hat{y})=V^{-1}_{\hat{y}}(y)$. Nevertheless, the derivatives of these two objects show subtle, but important differences.